# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ 5.3 Small ball probability
Small ball probability is directly linked to the $\epsilon$-entropy of the RKHSs (Lemma I.29 and I.30 of [22]), provides lower limits for empirical processes (Section 7.3 of [37]) and plays a crucial role in the convergence rate of the corresponding posterior distribution [52, 51]. For Brownian motion, the small ball probability can also be interpreted as the first exit time. For more details and applications, see for instance the survey [38] and references therein.
The following lemma provides a lower bound for small ball probability or equivalently an upper bound for the small ball exponent $\ln\mathrm{Pr}(\|\hat{Z}^{\tau,s}\|_{\infty,n}<\epsilon)$ of the Vecchia GP.
@@
Under Conditions 1 and 2, if the mother Gaussian process is Matérn with smoothness parameter $\alpha$, then for $\epsilon>0$ small enough,
@@
The proof of Theorem 2 follows a different route from the standard techniques applied for small deviation bounds of stationary Gaussian processes. Instead of deriving the small ball probability via $\epsilon$-entropy of the corresponding Reproducing Kernel Hilbert Space, we directly study it by utilizing the conditional distribution results derived in the previous section. The techniques are inspired by Chapter 7 of [33] and generalized in our proof. We believe these techniques are of independent interest and may be employed to analyze a broader class of stochastic processes.
For scaling parameters $\tau=s=1$, the lower bound for the small ball probability in Theorem 2 coincides with the Matérn process, see Lemma 11.36 of [22]. This indicates that despite having different covariance structures and lacking stationarity, the Vecchia approximation of a Matérn process has similar small deviation properties as the mother process. Such property is crucial in deriving matching, optimal posterior contraction rates for these two processes.
@@ 5.4 Decentered small ball probability
for $f_{0}\in L_{\infty}(\Omega)$. We start by recalling the definition of the RKHS associated to the Vecchia GP and discuss how its approximation properties relate to the decentered small ball probabilities.
@@
For all $f=\sum_{i=1}^{n}a_{i}\hat{K}^{\tau,s}(X_{i},\cdot)\in\mathbb{H}_{n}^{\tau,s}$, denote $a=(a_{1},a_{2},\cdots,a_{n})^{T}$, then the RKHS norm of $f$ takes the form
@@
We can now define the decentering function with argument $\epsilon>0$ as
$\inf_{f\in\mathbb{H}_{n}^{\tau,s}:\|f-f_{0}\|_{\infty,n}\leq\epsilon}\|f\|_{\mathbb{H}_{n}^{\tau,s}}^{2}.$
The combination of the decentering function and the centered small ball exponent results in the so called concentration function
@@
Finally, in view of Proposition 11.19 of [22], for any $f_{0}$ in the closure of the RKHS with respect to the empirical supremum norm $\|\cdot\|_{\infty,n}$, the decentered small ball probability satisfies that
@@
Therefore, understanding the asymptotic behaviour of the decentering function (as $\epsilon$ tends to zero with $n$) will provide us the decentered small ball probabilities.
To derive an upper bound for the decentering function we first provide some technical lemmas which are of interest on their own right. The first lemma discusses the minimal RKHS norm among all functions that interpolate a given set of points.
@@
Let $Z$ be a Gaussian process defined on $\mathcal{X}$ and $\mathbb{H}$ the corresponding RKHS. Then for any finite set $A=\{x_{1},x_{2},\cdots,x_{|A|}\}$ and function $f_{0}:\mathcal{X}\to\mathbb{R}$,
@@
Moreover, the minimum in the above equation is obtained by the function $f(\cdot)=\sum_{i=1}^{|A|}a_{i}K(x_{i},\cdot)$, with coefficients $a=(a_{1},a_{2},\cdots,a_{|A|})^{T}=K_{A,A}^{-1}\bm{f}_{0}(A)$.
@@
Then by combining the above two lemmas together with the upper bound for the variance $\mathrm{Var}\big{\{}Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]\big{\}}$, see Theorem 1, we get a more explicit control on the approximation error of the Gaussian process interpolation.
@@
Let $f_{0}$ belong to $C^{\beta}_{1}$ with $\beta\leq\alpha$ and let $Z^{\tau,s}$ be the rescaled Matérn process with smoothness parameter $\alpha$. Under Conditions 1 and 2, for all $X_{i}\in\mathcal{X}_{n}$ and its parent set $\mathrm{pa}(X_{i})$, we have
@@
Building on the above technical results, we can provide the following upper bound for the decentering function.
@@
Finally, combining the upper bounds derived for the (centered) small ball probability and the decentering function of the Vecchia GP in Theorem 2 and Lemma 11, respectively, results in an upper bound for the concentration function (37). This in turn, following from (38), provides a lower bound for the decentered small ball probability.
@@ 6 Bayesian nonparametrics
In this section, we provide posterior contraction rate guarantees for the Vecchia GP in context of the nonparametric regression model. However, we note, that given the general theory for Gaussian processes, see [52], these results can be easily extended to other nonparametric models, including density estimation and classification. With the help of the probabilistic properties developed in the previous sections, we present two main theoretical results. First we derive contraction rates of the Vecchia GP $\hat{Z}^{\tau,s}$ for fixed scaling hyper-paramaters $\tau$ and $s$ and show that optimal, oracle choices of these parameters (depending on the regularity parameter $\beta$ of the regression function $f_{0}$) lead to minimax optimal rates. In the second part of the section we consider the hierarchical Bayesian framework, where the scaling hyper-parameters $(\tau,s)$ of the GP are endowed with a hyper-prior distribution. We show that this fully Bayesian approach can adapt to the minimax rate without using any knowledge of the underlying true function. The proofs for Theorem 3 and 4 and Corollary 1 and 2 are deferred to Section D of the Supplementary Material.
@@ 6.1 Posterior contraction rates
There is a rich literature of posterior contraction rates for Gaussian processes based on the concentration function $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})$, defined in (37), see e.g. [52, 22]. In fact, by relating the metric used for testing the regression function $f_{0}$ to the norm of the Banach space in which the GP resides, it can be shown that the posterior contracts around the truth with the rate $\epsilon_{n}$ with respect of the testing metric, satisfying $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})\lesssim n\epsilon_{n}^{2}$. We formalize this result in context of the nonparametric regression model by considering the empirical $L_{\infty}$ norm as the Banach space norm and the empirical $L_{2}$ norm as the testing metric, which also determines the posterior contraction rate.
@@
Let us assume that $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})\lesssim n\epsilon_{n}^{2}$ holds for some $\epsilon_{n}\to 0$ for all $f_{0}\in C^{\beta}_{1}$. Then for every sequence of $M_{n}\to\infty$, the posterior on the regression function $f$ satisfies
@@
In the previous section we have derived lower bounds for the decentered small ball probability in Lemma 12 using the upper bound on the concentration function obtained by combining Theorem 2 and Lemma 11. Therefore, by plugging in this upper bound into the concentration inequality, one can obtain the contraction rate of the posterior in view of Lemma 13.
@@
Let $f_{0}\in C^{\beta}_{1}$, with $\beta\leq\alpha$, and $\hat{Z}^{\tau,s}$ be the the Vecchia approximation of the rescaled Matérn process with smoothness parameter $\alpha$. Under Conditions 1 - 2,
@@
Note, that the obtained rate in Theorem 3 crucially depends on the choice of the rescaling hyper-parameters $\tau$ and $s$ and the operator norm upper bound of the recursive GP interpolation $\tilde{\vartheta}_{n}$. We have shown that the latter is uniformly bounded in case of grid data. Assuming more generally its boundedness, we can omit it from the rate. Furthermore, if both $\tau$ and $s$ are constants, we retrieve the well known posterior contraction rates derived in [52]. This rate is minimax optimal if the regularity of the prior matches the regularity of the underlying true function, i.e. $\alpha=\beta$. At the same time, by allowing $s$ and $\tau$ to depend on $n$ and minimizing the right hand side of equation (3) one can obtain the minimax rate $n^{-\beta/(2\beta+d)}$ even in the case when the regularity of the prior and $f_{0}$ is not coinciding, see also [22]. We summarize these in the following corollary.
@@
Suppose that the conditions of Theorem 3 and $\tilde{\vartheta}=O(1)$ hold. If $\tau\asymp s\asymp 1$, then the posterior contracts with rate $\epsilon_{n}=n^{-\frac{\beta}{2\alpha+d}}(\ln n)^{\mathbbm{1}_{\{\alpha\in\mathbb{N}\}}}$, while for $\tau^{\alpha}s=n^{\frac{\alpha-\beta}{2\beta+d}}$ and $s\gtrsim 1$ one achieves the minimax contraction rate $\epsilon_{n}=n^{-\beta/(2\beta+d)}(\ln n)^{\mathbbm{1}_{\{\alpha\in\mathbb{N}\}}}$.
Theorem 3 and Corollary 1 show that the Vecchia approximations of Matérn processes retain exactly the same posterior contraction rates as the mother Matérn processes, see e.g. [48]. These verify that one can enjoy the benefits of scalable computation with Vecchia Gaussian processes without any loss in terms of estimation error rate. Hence, our results fill the gap for the statistical guarantees of Vecchia GPs that have been manifesting in the past few years.
@@ 6.2 Adaptation with hierarchical Bayes
While Corollary 1 recovers the optimal minimax contraction rate for arbitrary $\alpha\geq\beta$, this can be achieved only for some oracle choices of the scaling parameters $\tau,s$, depending on the smoothness $\beta$ of the true regression function $f_{0}$. This information, however, is typically not available in real world applications. Therefore, a natural idea is to introduce another layer of prior on the scaling parameters $\tau$ and $s$ such that they can automatically adapt to the dataset. Let us consider continuous priors on the hyper-parameters and denote the probability mass function of $(\tau,s)$ by $p(\tau,s)$. Then the hierarchical Vecchia GP $\hat{Z}_{x}^{\mathrm{Pr}}$ takes the form
@@
where $\hat{Z}^{\tau,s}$ denotes the Vecchia approximation of the rescaled Matérn process with smoothness parameter $\alpha$ and scale parameters $\tau$ and $s$. To formulate our contraction rate results, we need the following condition regarding the hyperprior on the scale parameters.
@@
The following theorem states our result regarding adaptation on $\tau$ and $s$.
@@
Let $f_{0}\in C^{\beta}_{1}$ for $\beta\leq\alpha$ and let $\hat{Z}_{x}^{\mathrm{Pr}}$ be the hierarchical Vecchia GP defined in (40) with hyper-prior satisfying Condition 3. If for all $\tau,s$ in the support of $p(\tau,s)$, the process $\hat{Z}^{\tau,s}$ satisfies Conditions 1 - 2 and the Gaussian interpolation bound $\tilde{\vartheta}_{n}=O(1)$, then for all $M_{n}\to\infty$,
@@
As a special case, one can fix either of the scaling hyper-parameters and endow the other with a hyper-prior. This would also lead to a rate adaptive procedure. The following corollary states this in case the time scaling hyper-parameter is set random and the other fixed.
@@
Let us fix $s=1$ and endow $\tau$ with a hyper-prior satisfying
@@
Then Condition 3 is satisfied and hence the posterior corresponding to the hierarchical Vecchia GP prior achieves the minimax contraction rate under Conditions 1 - 2 and $\tilde{\vartheta}_{n}=O(1)$.
@@
We note that adaptive contraction rate results for Matérn process has been derived in several papers. However, in most of these works the time scaling parameter was set to $\tau=1$ and adaptation was considered in the space scaling $s$. The only exceptions are the recent papers [48, 16]. In the former one the regularity $\alpha$ was set to half integers, while in the later the scale parameter $\tau$ was assumed to exceed 1. The major technical challenge is to characterize the RKHSs with different rescaling parameter $\tau$, which do not have a simple inclusion structure. With the help of the probabilistic properties developed in Section 5, we are able to prove the adaptation of Vecchia approximation of Matérn processes for arbitrary dimension and smoothness, and considering both the space and time rescaling hyper-parameters without the restrictions introduced in the previous papers.
7 Numerical analysis
The success of Vecchia Gaussian processes in practical applications have been widely demonstrated in the literature. In the numerical analysis we focus on validating the key mechanisms of Vecchia GPs that are conveyed throughout our theorems and proofs. Since Vecchia GP methods only differ in their DAG structures, we provide a universal package, called GPDAG, for posterior inference of Vecchia GPs to ensure fair comparison. This package is implemented in C++ using the state-of-the-art preconditioned conjugate gradient descent method [30] and comes with a high-level R interface. A detailed description of the posterior inference algorithms is available in Section E of the Supplementary Material.
7.1 Optimal estimation without prior approximation
In the literature, Vecchia GPs are often considered as approximations of their corresponding mother GPs. As a result, the goodness of a Vecchia GP is often measured by the approximation accuracy (in some proper metric) of its mother GP. In this section, we numerically demonstrate one of our key arguments, i.e. Vecchia GPs can perform optimal nonparametric estimation without providing a good approximation of the corresponding mother GP. Specifically, we consider the nonparametric regression problem and investigate the performance of the posterior resulting from the Vecchia GP prior (13). We implement Vecchia GPs with two different DAG structures. The first is the Layered Norming DAG (denoted as Norming in the figures) described in Section 4, see also Algorithm 1 for formalization. The second DAG structure is the Nearest Neighbor Gaussian Process [12] (NNGP) with maximin ordering [23] (denoted as Maximin in the figures). For simplicity, we set the interval $[0,1]$ as the domain and took the equidistant design on it as the training data $\mathcal{X}_{n}$, where $n$ ranges from $17$ to $2049$ (i.e. $n=2^{k}+1$, with $k$ ranging from 4 to 11). The true regression function is a randomly generated function from the sample paths of a Matérn process with smoothness $\alpha=3/2$ and scaling parameters $\tau=10,s=1$. We set the error standard deviation to $\sigma=0.1$ in equation (4). For both Vecchia GP methods we use the same Matérn process that generated the true $f_{0}$ above. In other words, the prior smoothness exactly matches the truth and the Vecchia GP with the Layered Norming DAG shall achieve the minimax rate according to our theory.
The DAG structures of the considered two Vecchia GP methods are as follows. For the Layered Norming DAG in Algorithm 1, since $\underline{\alpha}=1$ and $d=1$, the norming parent sets have cardinality $2$. This means for $X_{i}\in\mathcal{N}_{j}$, $\mathrm{pa}(X_{i})$ is the two closest locations to $X_{i}$ in $\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$. For NNGP with maximin ordering, we choose the neighborhood size as the closest integer to $2\ln(n)$. While there are disputes regarding the size of the parent sets in NNGP, previous works propose a logarithmic rate, see for instance [25, 44].
We illustrate the posterior contraction rate in the nonparametric regression model using Vecchia GPs in Figure 5. Both Vecchia GP methods show similar performances, with relatively large estimation error and wide credible bands when the sample size is small ($n=33$), but they can estimate the true function perfectly well as the sample size becomes large enough ($n=513$). Figure 6 provides a more qualitative report of the two Vecchia GPs in view of estimation accuracy, prior approximation and run time. The estimation error is measured by the $\ell^{2}$-distance between the posterior mean and the true regression function. As the sample size $n$ increases, both Vecchia GP methods recover the true regression function similarly well in $\ell^{2}$-norm. The prior approximation error is measured by the squared Wasserstein-2 distance between the marginal distributions of the Vecchia GP and the corresponding mother GP on $\mathcal{X}_{n}$. As one can observe, the Vecchia GP with Layered Norming DAG does not approximate the mother GP well. In fact, the squared Wasserstein-2 distance does not go below $20$ for any sample size $n$, while the Vecchia GP with maximin ordering has a squared Wasserstein-2 distance tending to zero. In view of the posterior estimation and prior approximation results, it is evident that Vecchia GPs do not require close approximation of the corresponding mother GPs, but it is sufficient to inherit some key regularity properties to recover the true function.
We also plot the computation times of the two Vecchia GPs on the right side of Figure 6. Since the Vecchia GP with Layered Norming DAG has fixed cardinality $2$ for all parent sets while NNGP with Maximin ordering has parent sets of size $2\ln(n)$, it is intuitively clear that the former can be computed much faster than the later. The difference in the runtime increases with the sample size. Therefore, if the objective is to recover the underlying true function, we can employ Vecchia GPs with very small parent set size ($2$ in this example), exploiting the computational benefits without loss of statistical efficiency.
Figure 5: Illustration of Nonparametric Regression with Vecchia GPs. Black lines and dots denote the true regression function and the observed data, respectively. Colored lines and regions denote the posterior mean and the $95\%$ pointwise credible intervals from the two Vecchia GP methods.
Figure 6: Qualitative results of the two Vecchia GP methods (Norming and Maxmin) when the true function has Hölder regularity $\beta=1.5$. Left: posterior estimation error measured by the $\ell^{2}$ distance between the true regression function $f_{0}$ and the posterior mean; Middle: prior approximation error measured by squared Wasserstein distance between marginals of the Vecchia GPs and their mother GPs; Right: Run time of MCMC inference measured in seconds.
7.2 Infinite smooth truth and finite norming sets
According to our theoretical results, choosing parent sets as norming sets with cardinality $\underline{\alpha}+d\choose\underline{\alpha}$ leads to optimal nonparametric estimation rates. While for Matérn processes with fixed regularity $\alpha$, such parent sets are easily implementable, for the Gaussian processes with squared exponential covariance kernel (which can be regarded as the limit of Matérn process with $\alpha\to\infty$), the corresponding cardinality would be infinite, hence practically infeasible. A natural question is, how to choose the parent sets for Gaussian processes with very high or even analytic regularity?
The rigorous theoretical investigation of this question is beyond the scope of the present paper. Nevertheless, we conjecture that if both the truth and the prior are $\alpha$ smooth for some large value $\alpha\in(0,+\infty]$, but we choose the parent sets as norming sets with cardinality $\tilde{\alpha}+d\choose\tilde{\alpha}$ for some integer $\tilde{\alpha}<\underline{\alpha}$, then the posterior will contract at a slower rate, depending on $\tilde{\alpha}$.
We demonstrate our conjecture through a numerical analysis. We take a sample path from a uni-dimensional Gaussian process with squared-exponential covariance kernel as a true function. Note, that this provides an analytic truth (with probability one). Then we compare three Vecchia GP methods. All methods use the squared exponential covariance kernel as the mother GP. So from the Bayesian perspective, the prior belief regarding the regularity of the truth is correct. Similarly to Section 7.1, these Vecchia GP methods only differ in their DAG structures. The method denoted as Maximin is the Nearest Neighbor Gaussian Process with maximin ordering, the same one used in the previous subsection with parent sets of size $2\ln(n)$. The method denoted as Norming2 and Norming4 employ the Layered Norming DAG with norming sets of cardinality 2 and 4, respectively.
The results of the numerical studies are shown in Figure 7. For the Maximin method, since the covariance matrix for the parent set becomes “numerically singular”, even using double precision floating numbers in C++, we could only conduct numerical studies up to sample size $n=257$. As expected, the Norming2 method provides a less accurate estimation of $f_{0}$ than Maximin. What is perhaps a bit surprising is that Norming4 provides a similarly good estimation as Maximin. This is a result of two phenomena. First, note that for $\alpha=\beta=4$ and $d=1$ the optimal posterior contraction rate is $n^{-\beta/(2\beta+d)}=n^{-4/9}$, very close to the minimax rate for estimating analytic functions $n^{-1/2}$, hence visually they are hard to distinguish. In this case $\underline{\alpha}=3$ hence the size of the norming set is ${\underline{\alpha}+d\choose\underline{\alpha}}=4$. Therefore, a Vecchia approximation with 4 parents results in a qualitatively similar estimation to the original squared exponential covariance kernel. Second, as we increase the size of the parent sets, the covariance matrix of the Vecchia GP starts to resemble more and more the covariance matrix of the mother GP. Therefore, the smallest eigenvalue of the Vecchia covariance matrix decays fast to zero, which in turn implies that the numerical error tends to dominate the statistical error.
Apart from the posterior estimation error, the prior approximation and the run time results behave as expected. The Layered Norming DAGs are not meant to approximate the mother process, hence not surprisingly the prior approximation error for Norming2 and Norming4 are large compared to the Maximin method. They do, however, take substantially less time and actually can be computed for larger sample sizes as well, compared to NNGP with maximin ordering.
Figure 7: Qualitative results of three Vecchia GP methods for analytic truth. Left: posterior estimation error measured by $\ell^{2}$ norm between the truth and the posterior mean; Middle: prior approximation error measured by squared Wasserstein distance between marginals of Vecchia GPs and their mother GPs; Right: Run time of MCMC inference measured in seconds.

### v4

@@ 4.3 Small ball probability
Small ball probability is directly linked to the $\epsilon$-entropy of the associated RKHSs (Lemma I.29 and I.30 of [22]), provides lower bounds for empirical processes (Section 7.3 of [38]) and plays a crucial role in the convergence rate of the corresponding posterior distribution [57, 56]. For Brownian motion, the small ball probability can also be interpreted as the first exit time. For more details and applications, see for instance the survey [39] and references therein.
The following lemma provides a lower bound for the small ball probability, or equivalently, an upper bound for the small ball exponent $\ln\mathrm{Pr}(\|\hat{Z}^{\tau,s}\|_{\infty,n}<\epsilon)$ of the Vecchia GP.
@@
Under Conditions 1 and 2, if the mother GP is a Matérn process with smoothness parameter $\alpha$, then for sufficiently small $\epsilon>0$,
@@
The proof of Theorem 2 follows a different route from the standard techniques used to obtain small deviation bounds of stationary Gaussian processes. Instead of deriving the small ball probability via the $\epsilon$-entropy of the corresponding Reproducing Kernel Hilbert Space, we study it by exploiting the conditional distribution results derived in the previous section. The techniques draw inspiration from Chapter 7 of [34] and are generalized in our analysis. We believe these techniques are of independent interest and may be applicable to a broader class of stochastic processes.
For scaling parameters $\tau=s=1$, the lower bound for the small ball probability in Theorem 2 coincides with that of the Matérn process, see Lemma 11.36 of [22]. This indicates that, despite having different covariance structures and lacking stationarity, the Vecchia approximation of a Matérn process exhibits similar small deviation behavior as the mother process. Such property is crucial for deriving matching, optimal posterior contraction rates for the two processes.
@@ 4.4 Decentered small ball probability
for $f_{0}\in L_{\infty}(\Omega)$. We begin by recalling the definition of the RKHS associated with the Vecchia GP and discuss how its approximation properties relate to decentered small ball probabilities.
@@
For any function of the form $f=\sum_{i=1}^{n}a_{i}\hat{K}^{\tau,s}(X_{i},\cdot)\in\mathbb{H}_{n}^{\tau,s}$, let $a=(a_{1},a_{2},\cdots,a_{n})^{T}$. Then the RKHS norm of $f$ is given by
@@
We can now define the decentering function as
$\epsilon\mapsto\inf_{f\in\mathbb{H}_{n}^{\tau,s}:\|f-f_{0}\|_{\infty,n}\leq\epsilon}\|f\|_{\mathbb{H}_{n}^{\tau,s}}^{2}.$
The combination of the decentering function and the centered small ball exponent yields the so called concentration function
@@
Finally, in view of Proposition 11.19 of [22], for any $f_{0}$ in the closure of the RKHS with respect to the empirical supremum norm $\|\cdot\|_{\infty,n}$, the decentered small ball probability satisfies
@@
Therefore, understanding the asymptotic behavior of the decentering function (as $\epsilon=\epsilon_{n}$ tends to zero with $n$) will provide us the decentered small ball probabilities.
To derive an upper bound for the decentering function, we first present several technical lemmas that are of independent interest. The first lemma discusses the minimal RKHS norm among all functions that interpolate a given set of points.
@@
Let $Z$ be a Gaussian process defined on $\mathcal{X}$ and $\mathbb{H}$ be the corresponding RKHS. Then, for any finite set $A=\{x_{1},x_{2},\cdots,x_{|A|}\}$ and any function $f_{0}:\mathcal{X}\to\mathbb{R}$,
@@
The minimum in the above equation is obtained by the function $f(\cdot)=\sum_{i=1}^{|A|}a_{i}K(x_{i},\cdot)$, with coefficients $a=(a_{1},a_{2},\cdots,a_{|A|})^{T}=K_{A,A}^{-1}\bm{f}_{0}(A)$.
@@
Then by combining the above two lemmas with the upper bound for the variance $\mathrm{Var}\big\{Z^{\tau,s}_{X_{i}}-\mathbb{E}[{Z}^{\tau,s}_{X_{i}}|{Z}^{\tau,s}_{\mathrm{pa}(X_{i})}]\big\}$, see Theorem 1, we obtain a more explicit control on the approximation error of the Gaussian process interpolation.
@@
Let $f_{0}\in C^{\beta}_{1}$ with $\beta\leq\alpha$, and let $Z^{\tau,s}$ be the rescaled Matérn process with smoothness parameter $\alpha$. Under Conditions 1 and 2, for all $X_{i}\in\mathcal{X}_{n}$ and its parent set $\mathrm{pa}(X_{i})$, we have
@@
Building on the technical results above, we can now provide the following upper bound for the decentering function.
@@
Finally, by combining the upper bounds derived for the (centered) small ball probability and the decentering function of the Vecchia GP in Theorem 2 and Lemma 8, respectively, we obtain an upper bound for the concentration function (32). This, in turn, via (33), provides a lower bound for the decentered small ball probability.
@@ 5 Bayesian nonparametrics
In this section, we provide posterior contraction rate guarantees for the Vecchia GP in the context of the nonparametric regression model. We note, however, that given the general theory for Gaussian processes (see [57]), these results can be readily extended to other nonparametric models, including density estimation and classification. Building on the probabilistic properties developed in the previous sections, we present two main theoretical results. First, we derive contraction rates for the Vecchia GP $\hat{Z}^{\tau,s}$ with fixed scaling hyperparameters $\tau$ and $s$, showing that optimal, oracle choices of these parameters – depending on the regularity parameter $\beta$ of the regression function $f_{0}$ – achieve minimax optimal rates. In the second part of the section, we consider a hierarchical Bayesian framework, where the scaling hyperparameters $(\tau,s)$ are endowed with a hyper-prior distribution. We show that this fully Bayesian approach can adapt to the minimax rate without requiring any knowledge of the underlying true function.
@@ 5.1 Posterior contraction rates
There is a rich literature on posterior contraction rates for Gaussian processes based on the concentration function $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})$, defined in (32), see e.g. [57, 22]. In fact, by relating the metric used for testing the regression function $f_{0}$ to the norm of the Banach space in which the GP resides, it can be shown that the posterior contracts around the truth at the rate $\epsilon_{n}$ with respect to the testing metric, provided that $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})\lesssim n\epsilon_{n}^{2}$. We formalize this result in the context of the nonparametric regression model by taking the empirical $L_{\infty}$ norm as the Banach space norm and the empirical $L_{2}$ norm as the testing metric, which in turn determines the posterior contraction rate.
@@
Assume that $\phi_{f_{0},n}^{\tau,s}(\epsilon_{n})\lesssim n\epsilon_{n}^{2}$ holds for some $\epsilon_{n}\to 0$ for all $f_{0}\in C^{\beta}_{1}$. Then, for every sequence $M_{n}\to\infty$, the posterior distribution of the regression function $f$ satisfies
@@
In the previous section we derived lower bounds for the decentered small ball probability in Lemma 9, using the upper bound on the concentration function obtained by combining Theorem 2 and Lemma 8. By plugging this upper bound into the concentration inequality, one can obtain the posterior contraction rate in view of Lemma 10.
@@
Let $f_{0}\in C^{\beta}_{1}$, with $\beta\leq\alpha$, and let $\hat{Z}^{\tau,s}$ be the Vecchia approximation of the rescaled Matérn process with smoothness parameter $\alpha$. Under Conditions 1 - 2,
@@
Note, that the rate obtained in Theorem 3 depends crucially on the choice of the rescaling hyperparameters $\tau$ and $s$, as well as on the operator norm upper bound of the recursive GP interpolation $\tilde{\vartheta}_{n}$. We have shown that $\tilde{\vartheta}_{n}$ is uniformly bounded in case of grid data. More generally, assuming its boundedness allows us to omit it from the rate. Furthermore, if both $\tau$ and $s$ are constants, we retrieve the well-known posterior contraction rates derived in [57]. This rate is minimax optimal when the prior’s regularity matches the regularity of the underlying true function, i.e. $\alpha=\beta$. At the same time, by allowing $s$ and $\tau$ to depend on $n$ and minimizing the right-hand side of equation (3) one can obtain the minimax rate $n^{-\beta/(2\beta+d)}$ even when the regularity of the prior does not match the regularity of $f_{0}$, see also [22]. These results are summarized in the following corollary.
@@
Suppose the conditions of Theorem 3 hold and $\tilde{\vartheta}=O(1)$. If $\tau\asymp s\asymp 1$, the posterior contracts at the rate $\epsilon_{n}=n^{-\frac{\beta}{2\alpha+d}}(\ln n)^{\mathbbm{1}_{\{\alpha\in\mathbb{N}\}}}$. On the other hand, if $\tau^{\alpha}s=n^{\frac{\alpha-\beta}{2\beta+d}}$ with $s\gtrsim 1$, one achieves the minimax contraction rate $\epsilon_{n}=n^{-\beta/(2\beta+d)}(\ln n)^{\mathbbm{1}_{\{\alpha\in\mathbb{N}\}}}$.
Theorem 3 and Corollary 1 show that the Vecchia approximations of Matérn processes achieve the same posterior contraction rates as their mother GP counterpart, see e.g. [51]. This confirms that one can enjoy the computational scalability of Vecchia Gaussian processes without any loss in estimation accuracy. In this way, our results address a long-standing gap in the statistical guarantees for Vecchia GPs that has persisted in recent years.
@@ 5.2 Adaptation with hierarchical Bayes
While Corollary 1 recovers the optimal minimax contraction rate for arbitrary $\alpha\geq\beta$, this can be achieved only for certain oracle choices of the scaling hyperparameters $\tau,s$, depending on the smoothness $\beta$ of the true regression function $f_{0}$. In practice, however, this information is typically unavailable. A natural solution is to introduce another layer of prior on the scaling parameters $\tau$ and $s$, allowing them to adapt automatically to the dataset. Let us consider continuous priors on these hyperparameters and denote the associated probability density function by $p(\tau,s)$. Then, the hierarchical Vecchia GP $\hat{Z}_{x}^{\mathrm{Pr}}$ takes the form
@@
where $\hat{Z}^{\tau,s}$ denotes the Vecchia approximation of the rescaled Matérn process with smoothness parameter $\alpha$ and scale parameters $\tau$ and $s$. To formulate our contraction rate results, we require the following condition on the hyperprior for the scale parameters.
@@
The following theorem presents our result regarding adaptation over $\tau$ and $s$.
@@
Let $f_{0}\in C^{\beta}_{1}$ with $\beta\leq\alpha$ and let $\hat{Z}_{x}^{\mathrm{Pr}}$ be the hierarchical Vecchia GP defined in (35) with hyper-prior satisfying Condition 3. If, for all $\tau,s$ in the support of $p(\tau,s)$, the process $\hat{Z}^{\tau,s}$ satisfies Conditions 1 - 2 and the Gaussian interpolation bound $\tilde{\vartheta}_{n}=O(1)$, then, for any sequence $M_{n}\to\infty$,
@@
As a special case, one can fix one of the scaling hyperparameters and endow the other with a hyper-prior. This also yields a rate-adaptive procedure. The following corollary illustrates this when the time-scaling hyperparameter $\tau$ is random and the other, $s$, is fixed.
@@
Fix $s=1$ and endow $\tau$ with a hyper-prior satisfying
@@
Then Condition 3 is satisfied and consequently the posterior corresponding to the hierarchical Vecchia GP prior achieves the minimax contraction rate under Conditions 1 - 2 and $\tilde{\vartheta}_{n}=O(1)$.
@@
We note that adaptive contraction rate results for Matérn processes have been derived in several papers. However, in most of these works, the time-scaling parameter was fixed at $\tau=1$ and adaptation was considered only with respect to the space-scaling parameter $s$. The only exceptions are the recent papers [51, 16]. In the former one the regularity $\alpha$ was restricted to half integers, while in the later one the time-scale parameter $\tau$ was assumed to exceed 1. The major technical challenge lies in characterizing the RKHSs corresponding to different rescaling parameters $\tau$, which do not possess a simple inclusion structure. By leveraging the probabilistic properties developed in Section 4, we are able to prove adaptation for Vecchia approximations of Matérn processes in arbitrary dimensions and smoothness levels, simultaneously considering both space and time rescaling hyperparameters without the restrictions imposed in previous works.